The evaporative condenser of an ammonia refrigeration plant has a water flow rate of 230 kg/s and enters the natural draft cooling tower at 40 C. The water is cooled to 30 C by air entering at 36 C db and 38% RH. The air leaves the tower as saturated at 45 C db. Calculate:

(a) Amount of makeup water in kg/hr.

(b) The cooling tower efficiency (%).


1
Expert's answer
2021-03-22T08:36:10-0400

a) Water make-up (M ) = Total water losses = Drift Losses ( D) + Evaporation Losses (E ) + Blow down Losses (B)

D = 0.1 to 0.3 percent of Circulating water (C ) for an induced draft cooling tower



C = Circulating water in m3/hr

 "\lambda" = Latent heat of vaporization of water = 540 kcal/kg or 2260 kJ / kg

Ti – T0 = water temperature difference from tower top to tower bottom in °C ( cooling tower inlet hot water and outlet cold water temperature difference)

Cp = specific heat of water = 1 kcal/kg / °C (or) 4.184 kJ / kg / °C

C = 230 kg/s = (230 kg / 998 kg/m3) / (1/3600 h) = 830 m3/h

E = (830 m3/h x (40 - 30°C) x 4.184 kJ / kg / °C ) / 2260 kJ / kg = 15.4 m3/h

M = 830 x 0.1 + 15.4 = 98.4 m3/h or 98 203 kg/h

b) Cooling tower efficiency can be expressed as

μ = (ti - to) 100 / (ti - twb)               

where

μ = cooling tower efficiency (%) - the common range is between 70 - 75%

ti = inlet temperature of water to the tower (oC)

to = outlet temperature of water from the tower (oC)

twb = wet bulb temperature of air (oC, oF)

μ = (40 - 30) / (40 - 26) * 100 = 71.4%